# Molecular Communication Channel as a Physical Reservoir Computer

### v1

@@ IV Results
In this section, we present the results of the performance evaluation for the MC reservoir computer using two complementary approaches. First, deterministic analyses were carried out in MATLAB using the mean‑field kinetic model (Eq. (6)) with the default simulation settings listed in Table I. Second, to capture the stochasticity of the MC channel, we performed particle‑based spatial stochastic simulations in Smoldyn [41]. By comparing the outcomes of the deterministic and stochastic simulations, we both confirm the validity of the approximate deterministic model and quantify the effects of intrinsic noise on RC performance.
@@
Mackey–Glass prediction length $\Delta$ $6$ steps
@@
Tx–Rx distance $x_{0}$ $10\,\upmu\mathrm{m}$
Reservoir nodes $M$ $100$
Numerical (deterministic) soln. timestep $\Delta t_{d}$ $1\,\mathrm{ms}$
Stochastic simulation timestep $\Delta t_{s}$ $10\,\mathrm{ms}$
Stochastic simulation receiver radius $R$ $3\,\upmu\mathrm{m}$
@@
IV-A Chaotic Time-Series Prediction
We benchmark our MC reservoir computer on two widely used chaotic time‑series prediction tasks: Mackey–Glass and NARMA10 [42]. These tasks have become standards for assessing both algorithmic and physical RC architectures because they combine complex temporal dependencies, strong nonlinearities, and a requirement for long‑range memory. The goal is to learn a mapping from the present (and past) states of the time series to future values, using only a linear readout layer trained on the reservoir states.
Performance Metric: We quantify the predictive accuracy of the reservoir using the normalized root-mean-square error (NRMSE),
where $y_{k}$ is the true target value at time $k$, $\hat{y}_{k}$ is the reservoir’s predicted value, $\bar{y}$ is the mean of all true target values, and $N_{\mathrm{test}}$ is the number of test samples. An NRMSE closer to 00 indicates a more accurate prediction.
Figure 4: Effects of MC reservoir parameters on computational performance in terms of NMRSE, (a-c) for Mackey-Glass series, and (d-f) for NARMA10 series prediction tasks.
IV-A1 Mackey-Glass Series
The Mackey–Glass system is described by a time-delayed differential equation originally introduced to model physiological control processes such as blood production:
@@ IV-A1 Mackey-Glass Series
where $\beta$, $\gamma$, $n$, and $\tau$ are positive constants governing the nonlinear feedback loop. For the canonical parameter set $\beta=0.2$, $\gamma=0.1$, $n=10$, and $\tau=17$, the system exhibits chaotic behavior with a fractal attractor.
We generate a univariate time series $\{x_{k}\}$ by numerically integrating the Mackey–Glass equation using a suitable step size (e.g., $0.1$), then sampling at discrete time points $k$. The reservoir input at time step $k$ is set to $u(k)=x_{k}$, and the task is to predict $x_{k+\Delta}$, i.e., the value of the series $\Delta$ steps into the future.
IV-A2 NARMA10 Series
A complementary benchmark is the NARMA10 (tenth‑order Non‑linear Auto‑Regressive Moving‑Average) series, defined by
@@ IV-A2 NARMA10 Series
where $y(k)$ is the output sequence and $u(k)$ is a scalar driving input (often i.i.d. uniform in [0, 0.5]00.5[0,\,0.5][ 0 , 0.5 ]). In this one-step-ahead prediction task, the reservoir receives $u(k)$ at time $k$ and must produce $\hat{y}(k+1)$ that approximates the true next value $y(k+1)$. The main challenge is that $y(k+1)$ depends on its previous ten outputs in a highly nonlinear manner, requiring the reservoir to capture temporal dependencies spanning at least ten steps.
IV-A3 Discussion
Fig. 3 summarizes the MC reservoir’s predictive performance on these tasks under different simulation conditions (rows) and time horizons (columns). The leftmost column illustrates the reservoir’s one-step-ahead forecasts for the NARMA10 series; the remaining three columns correspond to Mackey–Glass predictions at horizons of 1, 10, and 20 steps, respectively. Within each panel, the true time series (blue) is overlaid with the predicted values (red).
Rows (a–d) in Fig. 3 show the deterministic mean-field simulation results. Rows (e–h) show particle-based stochastic simulations that capture molecular noise. We observe that although the more realistic stochastic simulations introduce additional fluctuations in the reservoir node states, the overall predictive accuracy remains quite comparable to the deterministic case, indicating that the approximate mean-field analysis is valid and that the fluctuations do not severely degrade performance.
Focusing on the Mackey–Glass predictions (columns 2–4), we see that the reservoir accurately tracks the chaotic trajectory for modest horizons (e.g., $\Delta=1$ or $\Delta=10$), but it can still capture essential dynamics even at $\Delta=20$. Unsurprisingly, the error grows at larger horizons, yet the predictions remain qualitatively coherent with the true signal. This capability underscores the MC reservoir’s capacity for handling long-range memory in a strongly nonlinear system.
In the leftmost column, the reservoir’s one-step predictions for NARMA10 closely match the target sequence. Although NARMA10 typically requires the memory of up to ten past outputs, the reservoir’s nonlinear fading dynamics prove sufficient to encode that history and generate accurate forecasts.
IV-B Optimization of the Reservoir
IV-B1 Effect on MC System Parameters on Performance
The computational efficacy of the proposed molecular reservoir computer is intricately related to the biophysical parameters that govern the MC channel. Specifically, the dynamics of ligand-receptor binding, molecular propagation, and symbol encoding collectively define the temporal complexity and nonlinearity accessible to the reservoir. To systematically evaluate these dependencies, we conducted a parametric study involving some key variables: the ligand-receptor binding and unbinding rate parameters governing the receiver response {$k_{on}$, $k_{off}$}, the symbol interval and input amplitude governing transmitted signal{$N_{max}$, $T$}, and the propagation channel parameters such as diffusion coefficient (D) and transmitter-receiver distance $x_{0}$. The effects of changing these parameters in a physically plausible range are presented in Fig. 4 through the metric of NRMSE.
Our findings reveal a nonmonotonic relationship between parameter settings and predictive performance, as measured by the NRMSE. Specifically, the interplay between binding dynamics (i.e., the ratio $k_{off}$/$k_{on}$), inter-symbol duration, diffusion rate, and transmitter-receiver distance plays a critical role in determining the channel’s computational effectiveness. Low $k_{off}$/$k_{on}$ ratios, short inter-symbol durations, and high diffusion coefficients—especially when combined with shorter distances—lead to receiver saturation. In these scenarios, the occupation ratio of the receptor remains consistently high, nearing full saturation with minimal temporal variation. As a result, the system exhibits reduced dynamical richness and a diminished capacity to encode time-varying input features, which leads to an increase in NRMSE and degraded predictive accuracy. This effect is particularly evident in high-density ligand environments where excessive receptor occupation eliminates the temporal diversity necessary for reservoir computation.
Conversely, conditions characterized by high $k_{off}$/$k_{on}$ ratios or insufficient diffusion coefficients impede the ability of ligands to effectively bind and form receptor-ligand complexes. In such settings, the MC channel fails to facilitate reliable signal transmission, leading to sparse or inconsistent occupation patterns. This not only reduces the input-driven modulation of the reservoir state but also undermines the system’s ability to perform the nonlinear transformation essential for projecting input signals into a higher-dimensional computational space. Moreover, the consistent trends in prediction performance observed across both the NARMA10 and Mackey–Glass benchmarks reinforce the generalizability of these parameter–performance relationships, demonstrating that the interplay between binding kinetics and diffusion dynamics fundamentally governs the temporal richness and nonlinearity of the reservoir, and thus provides clear guidelines for optimizing MC‑based reservoir computers under diverse biochemical conditions.
@@ IV-B1 Effect on MC System Parameters on Performance
IV-B2 Post-Process Filtering
Due to the inherent stochasticity of molecular diffusion and receptor-ligand interactions, especially in low-concentration regimes, the raw output of the molecular reservoir exhibits noise-driven fluctuations, as observed in Fig. 2. These fluctuations, though biologically realistic, can obfuscate the underlying information-carrying trends. To address this, we applied a moving average filter to the received signal of receptor occupation $b(t)$, effectively attenuating high-frequency noise components while preserving meaningful temporal dynamics. The entire received signal sequence is filtered, the effects of which on the prediction task are shown in Fig. 3. The post-processed received signals are observed to display outputs similar to those of the estimation layer to deterministic simulations, enabling more robust performance during inference. The change in error with respect to the length of the filtering window is presented in Fig. 5, the filter improves performance until a certain point, after which the loss of information becomes too great and leads to an increase in error.

### v2

@@ IV Results
Here, we present a comprehensive performance evaluation for the MC reservoir computer. The analysis is structured to first demonstrate the MC reservoir’s core capabilities, then to investigate the factors that govern its computational power. We begin by benchmarking the performance on two standard chaotic time-series prediction tasks: the Mackey–Glass and NARMA10 series. We investigate the impact of inherent molecular noise on the performance and demonstrate how post-process low-pass filtering can be used to improve prediction accuracy. We then conduct a comprehensive parametric analysis to reveal how key MC system parameters, such as receptor kinetics, diffusion rate, and transmitter-receiver distance, can be tuned to optimize performance. This analysis is carried out through two complementary metrics: the task-specific NRMSE and the task-independent IPC.
Unless otherwise specified, all analyses utilize the default parameters listed in Table I, wherein the parameters defining the physical MC system are chosen to represent physiologically relevant values consistent with established literature in MC [21, 22].
@@
Transmitter-receiver distance $d$ $10\,\upmu\mathrm{m}$
Receiver radius $R$ $3\,\upmu\mathrm{m}$
Receptor unbinding constant $k_{\text{off}}$ $1\,\mathrm{s^{-1}}$
Reservoir virtual nodes $M$ $100$
Mackey–Glass prediction length $\Delta$ 66 steps
@@
Numerical simulation timestep $\Delta t_{\mathrm{d}}$ $1\,\mathrm{ms}$
Stochastic simulation timestep $\Delta t_{\mathrm{s}}$ $10\,\mathrm{ms}$
IV-A Performance on Benchmark Time-Series Prediction Tasks
We benchmark the MC reservoir computer on two widely adopted chaotic time‑series prediction tasks: Mackey–Glass and NARMA10 [44]. These benchmarks are standards for evaluating RC architectures as they demand a combination of complex temporal dependency modeling, strong nonlinearity, and significant memory depth.
IV-A1 Computing Tasks and Performance Metric
Mackey–Glass Series: The Mackey–Glass system is described by a time-delayed differential equation originally introduced to model physiological control processes such as blood production:
@@ IV-A1 Computing Tasks and Performance Metric
where $\beta$, $\gamma$, $n$, and $\tau$ are positive constants governing the nonlinear feedback loop. For the canonical parameter set $\beta=0.2$, $\gamma=0.1$, $n=10$, and $\tau=17$, the system exhibits chaotic behavior with a fractal attractor. We generate a univariate time series $\{x_{k}\}$ by numerically integrating the Mackey–Glass equation, then sampling at discrete time points $k$. The reservoir input at time step $k$ is set to $u(k)=x_{k}$, and the task is to predict $x_{k+\Delta}$, i.e., the value of the series $\Delta$ steps into the future.
NARMA10 Series: A complementary benchmark is the NARMA10 (tenth‑order Non‑linear Auto‑Regressive Moving‑Average) series, defined by the recursive equation:
@@
where $y(k)$ is the output sequence and $u(k)$ is a scalar driving input (often i.i.d. uniform in [0, 0.5][0,\,0.5]). In this one-step-ahead prediction task, the reservoir receives $u(k)$ at time $k$ and must produce $\hat{y}(k+1)$ that approximates the true next value $y(k+1)$. The main challenge is that $y(k+1)$ depends on its previous ten outputs in a highly nonlinear manner, requiring the reservoir to capture temporal dependencies spanning at least ten steps.
Figure 3: Prediction performance of the MC reservoir on two chaotic time-series benchmarks. Columns: (left) one–step prediction of the NARMA10 sequence; (center–left, center–right, right) prediction of the Mackey–Glass sequence at prediction horizons of 1, 10, and 20 symbols, respectively. Rows: (a–d) deterministic mean‑field analysis; (e–h) particle‑based stochastic simulation; (i–l) stochastic simulation post‑processed with a $2000\,\text{ms}$ moving‑average filter. In every panel the blue curve is the ground‑truth target and the red curve is the reservoir prediction.
Performance Metric: Predictive accuracy is quantified using NRMSE:
where $y_{k}$ is the true target value at time $k$, $\hat{y}_{k}$ is the reservoir’s predicted value, $\bar{y}$ is the mean of all true target values, and $N_{\mathrm{test}}$ is the number of test samples. Lower NRMSE indicates higher accuracy.
IV-A2 Time-Series Prediction Results
Fig. 3 summarizes the MC reservoir’s predictive performance on the NARMA10 and Mackey–Glass tasks across three simulation settings: a deterministic mean-field model (a-d), a particle-based stochastic model capturing molecular noise (e-h), and the stochastic model with a post-processing filter applied to the reservoir state (i-l).
The deterministic simulations (Fig. 3a-d) provide a performance baseline, revealing the computational potential of the idealized MC reservoir. The reservoir achieves high-level accuracy on the one-step-ahead Mackey–Glass prediction (MG-1, NRMSE = 0.054), demonstrating its ability to capture short-term temporal dependencies. For longer horizons, we observe a non-trivial trend: performance degrades for MG-10 (NRMSE = 0.354) but then counter-intuitively improves for MG-20 (NRMSE = 0.311). We attribute this improvement not to a simple monotonic decay of memory, but to the specific autocorrelation function (ACF) of the Mackey–Glass series for $\tau=17$. The ACF for this chaotic series is quasi-periodic, exhibiting decaying oscillations (see, e.g., Fig. 2(b) in [41]). Our results suggest that for the given sampling rate, the prediction horizon of $\Delta=10$ steps corresponds to a lag time where the ACF magnitude is near a local minimum, i.e., close to a zero-crossing. At such a point, the linear correlation between the present state and the future target is weakest, making the prediction task inherently challenging for a linear readout layer and resulting in a higher NRMSE. In contrast, the $\Delta=20$ horizon appears to align with a region of resurgent correlation, i.e., a secondary peak in the ACF’s magnitude. Although this correlation is weaker than at very short lags, its greater magnitude compared to the lag at $\Delta=10$ provides the linear readout with a more stable and informative target, leading to a lower NRMSE. Therefore, the performance of the MC-PRC on chaotic time-series prediction is modulated not just by the reservoir’s memory decay, but also by the intrinsic correlation structure of the time-series itself.
The introduction of molecular noise in the particle-based stochastic simulations (Fig. 3e-h) expectedly degrades performance, yet reveals more complex dynamics such that the one-step-ahead Mackey–Glass prediction (MG-1) becomes significantly less accurate than the 20-step-ahead prediction (NRMSE of 0.865 vs. 0.702, respectively). We propose this counter-intuitive result stems from the relationship between the amplitude of the signal to be predicted and the inherent noise floor of the physical reservoir. For the MG-1 task, the target represents a subtle, incremental change in the smoothly-sampled time series. The stochastic fluctuations in the reservoir state, driven by discrete molecular events, are of a comparable magnitude and can therefore obscure this fine-grained signal from the linear readout. In contrast, the MG-20 prediction task targets a much larger-amplitude change due to the system’s chaotic divergence over a longer horizon. This larger signal remains clearly discernible above the reservoir’s intrinsic noise floor, allowing the readout to learn a more robust mapping and achieve a lower error.
To mitigate the detrimental effects of this noise, we apply a $2000$ ms moving-average filter to the reservoir’s raw output state before the readout stage (Fig. 3i-l). This post-processing step yields substantial improvements in predictive accuracy across all stochastic scenarios, with NRMSE values approaching those of the deterministic case (e.g., for MG-10, improving from $0.921$ to $0.583$). The quantitative effect of the filter window length, detailed in Fig. 4, reveals that the optimal filtering strategy is highly task-dependent. For the Mackey–Glass series, the NRMSE decreases monotonically with increasing window length before plateauing, suggesting its useful signal is encoded in low-frequency dynamics that benefit from aggressive noise removal. In stark contrast, the NARMA10 series exhibits a distinct U-shaped performance curve with an optimal window around $1000$ ms, indicating that essential, higher-frequency information would be erased by an overly long filter. Interestingly, this filtering alters the performance trend for Mackey–Glass predictions visible in Fig. 3: the MG-20 task (NRMSE = $0.653$) is now slightly less accurate than the MG-10 task (NRMSE = $0.583$). This reversal suggests that while the filter effectively removes detrimental high-frequency noise, its smoothing effect may also attenuate the subtle, low-frequency periodicities in the reservoir state that the unfiltered system exploited for the improved 20-step-ahead prediction.
@@ IV-A2 Time-Series Prediction Results
IV-B Parametric Analysis for Reservoir Optimization
We now conduct a comprehensive parametric analysis to understand the physical principles governing the MC reservoir’s performance and to identify its optimal operating regimes. We evaluate how key biophysical parameters influence the reservoir’s computational power using our two complementary metrics: the task-specific NRMSE and the task-independent IPC. To render this extensive parameter exploration computationally tractable, all subsequent analyses in this section are performed using the deterministic mean-field model.
Figure 5: Effects of MC reservoir parameters on computational performance in terms of NMRSE, (a-c) for Mackey–Glass series, and (d-f) for NARMA10 series prediction tasks.
IV-B1 Effect of MC System Parameters on Task Performance
The computational efficacy of the proposed MC reservoir is intricately governed by the biophysical parameters of the underlying communication channel. To elucidate these relationships, we systematically evaluated how performance on our benchmark tasks, quantified by NRMSE, depends on key physical parameters. The results, presented in Fig. 5, reveal clear operational regimes and trade-offs that are fundamental to designing such systems.
Receiver Kinetics
Figs. 5(a) and (d) explore the interplay between the receptor binding ($k_{\mathrm{on}}$) and unbinding ($k_{\mathrm{off}}$) constants. For both the Mackey–Glass and NARMA10 tasks, optimal performance (lowest NRMSE) is confined to a distinct diagonal region. This reveals that performance is not dictated by either rate alone, but by their ratio, which determines the receptor’s binding affinity (inversely related to the dissociation constant, $K_{\mathrm{D}}=k_{\mathrm{off}}/k_{\mathrm{on}}$). Two failure regimes are evident. At very low $k_{\mathrm{on}}$ or high $k_{\mathrm{off}}$ (low affinity), the system enters a sparse-binding regime where too few molecules are captured to form a reliable signal. Conversely, while a very low $k_{\mathrm{off}}$ (high affinity) indeed increases the reservoir’s memory by extending receptor occupancy time, this comes at the cost of dynamical responsiveness. In this regime, the reservoir state becomes dominated by its own long-timescale memory, effectively locking in past information. The influence of new, incoming molecular signals is consequently suppressed, as they are unable to significantly modulate the already-occupied receptor population. This loss of sensitivity to the input signal diminishes the reservoir’s ability to perform the nonlinear transformations necessary for computation. Therefore, optimal performance requires a finely tuned intermediate binding affinity. This balance provides sufficient memory to capture relevant temporal dependencies while preserving the crucial ability for the reservoir state to be dynamically and richly modulated by the present input.
Signal Encoding and Transmission Rate
Figs. 5(b) and (e) reveal that the optimal configuration of the number of released molecules ($N_{\mathrm{max}}$) and symbol duration ($T$) is fundamentally task-dependent. In our system, both parameters co-determine the reservoir’s two essential computational ingredients: fading memory and nonlinearity. The symbol duration $T$ directly tunes the structure of the memory by controlling ISI, while both $T$ and $N_{\mathrm{max}}$ jointly determine the effective molecular concentration at the receiver, which dictates the operating point on its nonlinear response curve. For the Mackey–Glass task (Fig. 5(b)), which requires predicting the evolution of a smooth, autonomous trajectory, the optimal region exhibits a diagonal trend where increasing $T$ requires a compensatory increase in $N_{\mathrm{max}}$. This suggests a need to maintain a consistent level of nonlinear activation; as pulses become more temporally isolated, they must be stronger to induce a comparable dynamical response. In contrast, the NARMA10 task (Fig. 5(e)) shows a markedly different behavior where performance is more strongly dependent on $N_{\mathrm{max}}$ than on $T$. Across all symbol durations beyond the initial high-ISI regime ($T>1$ s), performance consistently improves as $N_{\mathrm{max}}$ is decreased. We attribute this to the specific computational demands of the NARMA10 task, which, unlike the trajectory-following MG task, requires the nonlinear processing of a sequence of distinct, random input values.
Channel Propagation
Figs. 5(c) and (f) map performance across the diffusion coefficient ($D$) and transmitter-receiver distance ($d$), which together sculpt the channel’s impulse response and govern the signal strength that drives the receiver’s nonlinearity. The results reveal a striking diagonal band of optimal performance, demonstrating that computational success hinges not on independently tuning these parameters, but on achieving a critical balance between the memory timescale and the strength of nonlinear activation. In the regime of low $D$ and high $d$, severe signal attenuation weakens the arriving molecular pulse, causing the receiver to operate in a near-linear, computationally poor regime despite its long memory. Conversely, at high $D$ and low $d$, the characteristic transport time, scaling with $d^{2}/D$, becomes vanishingly short. The channel approaches a memoryless limit, and although the receiver’s nonlinearity is strongly engaged, it lacks the temporal context necessary for prediction. The optimal diagonal thus represents the sweet spot where the channel’s memory kernel is appropriately structured for the task, while the signal remains potent enough to drive the system into a computationally rich, nonlinear regime.
IV-B2 Information Processing Capacity (IPC) Analysis
To evaluate the computational capability of the MC reservoir more fundamentally, we employ the IPC framework introduced by Dambre et al. [45], and further refined by Schulte-to-Brinke et al. [46]. IPC quantifies how effectively the system states can approximate various nonlinear and memory-dependent transformations of past inputs, providing a comprehensive, task-independent measure of computational richness. Formally, IPC is computed by projecting the input history onto an orthonormal polynomial basis—specifically, the set of Legendre polynomials $\{P_{n}(\cdot)\}$, as they provide a complete orthonormal basis for signals defined on the interval [1,1][-1,1]. Given an input signal $u(t)$, normalized to have zero mean and constrained within the interval [1,1][-1,1], we define polynomial functionals of the form:
$y(t)=\prod_{i}P_{n_{i}}\left(u(t-s_{i})\right),$ (14)
where $n_{i}$ is the polynomial order and $s_{i}$ is the delay associated with the $i^{\text{th}}$ factor. Each unique set $\{(n_{i},s_{i})\}$ is termed a factor set and represents a specific computational task defined by polynomial order and memory depth.
We systematically enumerate factor sets using an ordered backtracking algorithm to ensure uniqueness and completeness while avoiding redundant evaluations [46]. For each factor set, we compute the optimal linear estimator $\hat{y}(t)$ from the system states $\mathbf{X}(t)$ via ridge regression (using pseudo-inverse methods) to avoid numerical instability. The corresponding capacity $C$ for each factor set is then defined as the squared correlation coefficient between the actual target $y(t)$ and its linear estimator $\hat{y}(t)$:
$C=\text{Corr}^{2}\left(y(t),\hat{y}(t)\right).$ (15)
Figure 6: Comparison of IPC contributions for different order and delay of univariate Legendre terms.
Figure 7: Parametric analysis of the MC reservoir’s IPC. The color in each panel represents the total IPC, summed over all significant first- and second-order Legendre polynomial terms. (a) IPC as a function of the receiver kinetics, governed by the binding ($k_{\mathrm{on}}$) and unbinding ($k_{\mathrm{off}}$) constants. (b) IPC as a function of the signal encoding, determined by the maximum number of released molecules ($N_{\mathrm{max}}$) and the symbol duration ($T$). (c) IPC as a function of the channel propagation properties, defined by the diffusion coefficient ($D$) and the transmitter-receiver distance.
Due to finite-sample estimation effects, very small IPC values may occur purely by chance. To mitigate these spurious capacities, we employ a statistical significance threshold based on the $\chi^{2}$-distribution as detailed in [45, 46]. Specifically, we select a significance threshold corresponding to a false-positive rate $p=10^{-4}$, and all IPC values below this threshold are set to zero. Total IPC is then the sum of all significant individual capacities:
$\text{IPC}_{\text{total}}=\sum_{\text{significant }C}C,$ (16)
thereby providing a rigorous measure of the system’s overall computational power. Finally, IPC results are presented either as total capacities, or subdivided by polynomial order, delay, or specific factor sets, facilitating a detailed interpretation of system performance.
Our analysis of IPC uncovers meaningful trends in how both polynomial order and memory delay influence capacity within the univariate Legendre polynomial basis (Fig. 6). As expected for any RC system, increasing the memory delay leads to a systematic reduction in IPC contribution. This reflects the inherent challenge of precisely reconstructing information from the distant past in a physical system, where signals naturally decay and become increasingly corrupted by temporal dispersion. Likewise, higher-order polynomials consistently yield lower IPC contributions. This is because more complex nonlinear transformations demand finer, more precise mappings from the reservoir state, which are inherently more difficult for a physical system with limited computational power to realize. Given the prominent roles played by first and second-order polynomial terms in relevant computational tasks, such as NARMA10 and Mackey–Glass time series prediction, our subsequent parametric analysis in Fig. 7 specifically focuses on the combined IPC of these orders.
Fig. 7(a) illustrates the effect of ligand–receptor interaction rates ($k_{\mathrm{on}}$ and $k_{\mathrm{off}}$) on IPC. The analysis reveals a well-defined optimal region at intermediate rates, surrounded by three distinct low-IPC regimes that correspond to the failure to achieve the necessary balance between memory and nonlinearity. We identify the following computationally poor regimes. First, at very low $k_{\mathrm{on}}$ and very high $k_{\mathrm{off}}$, the system is both memoryless and weakly coupled to the input; ligands bind rarely and unbind almost instantly, preventing the reservoir from either retaining past information or responding strongly to the present input. Second, in the regime of very low $k_{\mathrm{on}}$ and very low $k_{\mathrm{off}}$, the reservoir is effectively inert; the weak input coupling is combined with a static memory, meaning the state, once changed, fails to evolve in response to new information. The third and most critical failure mode occurs at very high $k_{\mathrm{on}}$ and very low $k_{\mathrm{off}}$. Here, the MC system enters deep receptor saturation. The high binding rate and long occupancy time cause the reservoir state to lock in, losing its dynamic range and sensitivity to input fluctuations. While this configuration provides a long memory, it is a static, computationally poor memory that suppresses the rich, graded nonlinear transformations essential for high IPC. The optimal region is therefore the sweet spot that successfully balances these competing factors.
Fig. 7(b) plots the IPC across the number of released molecules ($N_{\mathrm{max}}$) and the symbol duration ($T$). Consistent with our task-specific NRMSE analysis, the IPC is severely degraded at very short T due to the formation of unstructured memory, where catastrophic ISI destroys the temporal resolution necessary for computation. Beyond this initial high-ISI regime, IPC becomes largely independent of $T$ and is instead governed mainly by $N_{\mathrm{max}}$. This behavior is reminiscent of our findings for the NARMA10 task, suggesting that IPC (constructed with first and second-order terms) is more informative of the prediction performance in NARMA10 task compared to MG task.
Fig. 7(c) illustrates the IPC across the diffusion coefficient ($D$) and transmitter-receiver distance ($d$). A quasi-diagonal band of optimal IPC is observed, highlighting that computational richness is achieved by tuning the channel’s physical properties to match a characteristic timescale. In the region of low $D$ and high $d$, IPC is low. Here, signal attenuation is severe, leading to a weak signal at the receiver. This fails to activate the receiver’s nonlinearity, causing the system to operate in a near-linear, computationally poor regime. Although the memory kernel is long, its quality is poor due to high temporal dispersion and low signal strength.
These findings reveal the necessity of jointly tuning both physical and temporal system parameters to fully exploit the computational capacity of MC-based reservoir. The trends observed in the IPC results demonstrate a meaningful relationship with the reservoir’s empirical performance on time series prediction. Although IPC and NRMSE are distinct metrics, i.e., one measuring fundamental capacity, the other task-specific performance, their close behaviors across parameter sweeps make IPC a valuable theoretical tool for anticipating favorable operating conditions and guiding the design and tuning of MC-based reservoir computers. However, the precise quantitative link between a high total IPC and high performance on a specific task remains an open and challenging research question in the reservoir computing field [47].